if ray EG=59, what is the measure of EF i will try to illustrate it
E._____________F______________G
8x-14 4x+1
what does EF=

Answer by stanbon(75887)

if ray EG=59, what is the measure of EF i will try to illustrate it
E._____________F______________G
8x-14 4x+1
what does EF=
------
Since EG = 59, (8x-14)+(4x+1) = 59
12x - 13 = 59
12x = 72
x = 6
-----
Ans: EF = 8x-14 = 8*6-14 = 48-14 = 34
--------------------
Cheers,
Stan H.
----------